three coins are flipped 32 times. over these 32 trials, how many times would you expect to flip two heads and one tails.
Answer #1

Three coins are flipped together 32 times. In a single event, out of the 32 events, there are 8 possible outcomes (23=8) and so the probability of getting two heads and one tail is P = (2/8) = 0.25 = 1/4

so, 1 out of 4 times, one gets two heads and one tail in a single trial.

Since these coins are tossed 32 times ( 4 x 8), therefore, one must expect two heads and one tail 8 times in these 32 trials.